OpenStudy (anonymous):

The measures of the three angles of a triangle are represented by x,3x, and x+30 . Find the value of x . 1. 15 2. 30 c. 42 d. 65 Are they suppose to equal to 180 degrees ?

OpenStudy (anonymous):

Yep, total angles = 180

OpenStudy (mertsj):

Yes. Add them up and set the sum equal to 180. Solve for x

OpenStudy (anonymous):

so it would 5x+30=180 @Mertsj

OpenStudy (anonymous):

x+30+3x+x=180 5x+30=180 5x=150 x=30

OpenStudy (anonymous):

thanks both @Luis_Rivera @Mertsj

OpenStudy (mertsj):

yes, you are correct.